OpenStudy (anonymous):

How many solutions does the following equation have? 3x + 6 = 3(x + 2)

OpenStudy (anonymous):

factor the 3 on the right side and solve for x 3x+6 = 3x +6 everything cancels out and your left with no solution

OpenStudy (anonymous):

. A. exactly 2 solutions B. exactly 1 solution C. infinitely many solutions D. no solutions Those are the choices

OpenStudy (anonymous):

infinitely many

OpenStudy (anonymous):

So, its no solution?

OpenStudy (anonymous):

you can plug in any number and the equation will equal

OpenStudy (anonymous):

yea 3x - 3x = 0 so your left without a variable therefore its no solution

OpenStudy (anonymous):

How many do you want?

OpenStudy (anonymous):

Im confused.. gettin' three different answers.

OpenStudy (anonymous):

What do you mean Cliffsedge?

OpenStudy (anonymous):

Plug in any number you want for x and see if it works.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

@Kashan and @jayz657 are both incorrect.

OpenStudy (anonymous):

@jayz657 and @Kashan thats not right. the equation has infinitely many solutions. you can plug in infinitely many numbers and always get 0=0, which is true

jimthompson5910 (jim_thompson5910):

It can't be "no solution". Here's a counter-example showing why 3x + 6 = 3(x + 2) 3*0 + 6 = 3(0 + 2) ... Plug in x = 0 0 + 6 = 3(2) 6 = 6 .. which is TRUE So clearly x = 0 works and it is a solution

OpenStudy (anonymous):

Logic FTW.

OpenStudy (anonymous):

Ohh, so it is infinity ∞ ?

OpenStudy (anonymous):

X may be any real number as a solution.

OpenStudy (anonymous):

Ok thanks guys.